OpenStudy (anonymous):

Simplify square root of 2 multiplied by the cube root of 2.

Atsie (atsie):

Firstly, what is the square root of two?

Atsie (atsie):

Do it on the calculator, if you need to

OpenStudy (anonymous):

@Atsie 1.4

OpenStudy (tkhunny):

Have you considered fractional exponents?

OpenStudy (anonymous):

@tkhunny how would I do that?

OpenStudy (anonymous):

the X root of something is literally taking that number to the power of 1/X

OpenStudy (tkhunny):

You've never seen it? \(\sqrt{2} = 2^{1/2}\) \(\sqrt[3]{2} = 2^{1/3}\)

OpenStudy (anonymous):

@tkhunny no I haven't. then do i find the common denominator and add the exponents?

OpenStudy (tkhunny):

Demonstrate.

OpenStudy (anonymous):

@tkhunny 2^1/2 = 2^3/6 2^1/3 = 2^2/6 2^5/6

OpenStudy (tkhunny):

Does that look "simplified"?

OpenStudy (anonymous):

@tkhunny yes.

OpenStudy (tkhunny):

Well, there you go. Move on to the next one.

OpenStudy (anonymous):

@tkhunny thanks!